OpenStudy (anonymous):

HELP HELP HELP Will give medals and fan. square root of 8 x plus 1 equals 5 square root of x minus 7 end root plus 5 equals 11 negative 4 square root of x plus 9 equals 20 square root of 6 x plus 5 equals square root of 53 square root of x plus 2 end rot plus 4 equals x

OpenStudy (jdoe0001):

lemme try the 1st one, ok, see if you can do the others \(\bf \sqrt{8x+1} = 5\\ \textit{squaring both sides}\\ (\sqrt{8x+1})^2 = 5^2 \implies \sqrt{(8x+1)^2} = 5^2 \implies 8x+1 = 25\\ \textit{subtracting 1 from both sides}\\ 8x\cancel{+1(-1)} = 25(-1) \implies 8x = 24\\ \textit{dividing both sides by 8}\\ \cfrac{\cancel{8}x}{\cancel{8}} = \cfrac{24}{8} \implies x = \square?\)